OpenStudy (anonymous):

Multiply (2square root 5 - 4 square root 7)(4 square root 5 + 8 square root 7)=

OpenStudy (anonymous):

\[(2\sqrt{5}-4\sqrt{7})(4\sqrt{5}+8\sqrt{7})\]

OpenStudy (anonymous):

-184